# Fairly Accurate: Fairness-aware Multi-group Target Detection in Online Discussion

### v2

@@ 8 Results
Results show that $GAP_{multi}$ not only achieves better group-fairness by minimizing performance disparities impact across groups, but also maintains competitive overall accuracy. This highlights the effectiveness of $GAP_{multi}$ in addressing the unique challenges of multi-group target detection. All reported results in the main material are for the 7-group setting. For group performance on selected $\prescript{G\mkern-0.5mu}{}{C}_{2}$ isolated pairs, please refer to Appendix C.
8.1 Evaluation & Loss Performance
The values presented in Table 3 illustrate the achieved BA values across various groups over five runs for the test set, for the two baseline losses (OE and CLA) vs. $GAP_{multi}$.
Balanced Accuracy (BA) ($\uparrow$)
@@ 8.1 Evaluation & Loss Performance
OE 80.31 86.91 81.07 84.87 64.99 67.91 75.01 21.9$\pm$1.3 77.29$\pm$0.29 2/7
CLA 82.51 85.34 81.67 84.62 73.91 74.92 80.44 11.4$\pm$0.8 80.49$\pm$0.14 0/7
$GAP_{multi}$ 83.18 83.86 83.47 83.42 78.95 78.32 82.58 5.5$\pm$0.5 81.97$\pm$0.13 5/7
Table 3: Balanced Accuracy (BA) achieved by each loss function (OE and CLA baselines vs. $GAP_{multi}$) over the 7 demographic groups (on test data) for five runs. For each loss, we also color which group exhibits the maximum and minimum BA values achieved for that loss over the 7 groups, with the difference between this maximum vs. minimum shown in the Max. Diff. column (for BA, we want to minimize this difference). $GAP_{multi}$ achieves a lower maximum difference (Max. Diff. = 5.54) than either baseline loss functions, evident from the visualization in Fig. 3. $GAP_{multi}$ also achieves the highest (macro) average BA across groups (Avg. BA = 81.97), and the best BA for most of the groups (#Best BA column = 5/7). We report the standard deviation of Max. Diff and Avg. BA over five runs, with CLA and $GAP_{multi}$ having similar deviation. $GAP_{multi}$ achieves the best fairness performance while retaining overall utility.
Recall that Overall Error (OE) corresponds to minimizing overall error without any fairness constraints, and CLA equalizes false negative rates. The Black group, which constitutes the statistical majority group in the dataset (as shown in Figure 1), exhibits the maximum BA values for the three losses, with relatively similar performance across all three. Meanwhile, performance for the other groups, and in particular for the Native American group, which is a statistical minority, varies starkly across the different optimization strategies. Moreover, the performance for the group that exhibits the minimum BA values for each loss shows that the worse-off group exhibits the best performance under $GAP_{multi}$.
Figure 3: Visualization of the BA values achieved by each loss over the 7 demographic groups. The maximum difference (Max. Diff.) between the maximum and minimum BA achieved for each loss across groups is also shown. See Table 3 for additional details. $GAP_{multi}$ performs best with lowest Max Diff. of 5.5, showcasing that it greatly reduces the disparities in performance across groups vs. other losses.
We show the maximum difference (Max. Diff.) between the groups in Table 3, which captures the gap in performance between worst vs. best performing groups. We see that solely optimizing for overall error (OE) fails to account for variations in group performance, resulting in the highest difference values (Max. Diff. of 21.9). In contrast, both the $GAP_{multi}$ and CLA approaches incorporate considerations of group performance alongside overall error, leading to substantially lower Max. Diff. compared to OE. Notably, $GAP_{multi}$ optimizes for balanced error rates across groups, exhibiting the smallest difference (Max. Diff. of 5.5).
We report the average BA (macro) score obtained in all three losses in Table 3, with $GAP_{multi}$ having the highest value (Avg. BA: 81.97). We hypothesize that by incorporating the group information in the loss during model training, both the group-informed losses (CLA and $GAP_{multi}$) are able to find better local optima compared to OE. The Best BA is a group measure indicating how many amongst the 7 groups are performing best across the three losses. For this measure, we observe OE to have the highest group BA values for 2 groups (Black, Middle Eastern), while $GAP_{multi}$ performs best for rest of the five groups. This also emphasizes the fact that $GAP_{multi}$ does not prioritize the performance of one group over others during optimization, thereby being the best performing loss for most of the groups.
Figure 4: Heatmap of pairwise absolute difference of BA across groups in test set as an indicator for bias and disparate impact. OE has the highest performance gap (Max Diff = 21.9) across groups as indicated by the extremes of color, not only across one group- pair but consistently across multiple group pairs. $GAP_{multi}$ has the least spread in pairwise error values (Max Diff = 5.5), evident from the flatness of color, indicating least disparate impact across groups.
@@
The results for OE show that the (Black, Native American) pair has the largest Max. Diff. of 21.9. We also see other group pairs with wide performance disparity. Two key observations can be drawn. First, the Black group being the statistical majority group has a dominant performance gap over the Native American group, one of the statistical minorities. Second, even for group pairs that have similar prevalence in the data, e.g., the (Middle Eastern, White) pair, there can be performance disparities (BA gap of 9.9) because one group may simply be easier to identify than the other.
CLA, which optimizes for FNR, has improved performance over OE, however, since its optimization criteria (minimizing false negatives) does not align with the intended evaluation (having similar performance across groups), and its heatmap lies in-between that of OE and $GAP_{multi}$. In contrast, the $GAP_{multi}$ loss, which is explicitly designed to achieve similar (balanced) performance across groups while optimizing overall utility, shows substantially fewer extremes in group performance gaps. The heatmap reveals smoother transitions between groups, indicating equitable distribution of the classifier’s performance. Note that the extreme values of Maximum Difference (Max. Diff.) presented in Table 3 are reflected as outliers in these heatmaps, corresponding to specific group pairs.
OE CLA $GAP_{multi}$
Hamming Loss (HL) % ($\downarrow$) 7.65 12.10 6.85
Table 4: HL values across different losses for test data. Lower values are better. $GAP_{multi}$ optimizes for jointly over TPR and TNR, thereby achieving lowest values, indicating better classifier performance.
Hamming Loss (HL) (Eq. 11) values are shown in Table 4 as a summary statistic in our multi-label classification setup. The HL metric quantifies the average fraction of misclassified labels, with lower HL values indicating enhanced classifier performance. While the CLA loss (HL: 12.10) prioritizes minimizing FNR, which inherently involves asymmetry, there may be instances where it disproportionately optimizes for this aspect at the expense of other performance metrics. Consequently, CLA may exhibit poorer HL performance compared to OE loss (HL: 7.65). In contrast, the $GAP_{multi}$ loss maintains symmetry, resulting in a balanced trade-off while jointly minimizing TPR and TNR. As a result, $GAP_{multi}$ consistently achieves the lowest (HL: 6.85) values, indicative of superior classifier performance.
Loss $Prc_{macro}$ ($\uparrow$) $Rec_{macro}$ ($\uparrow$) $F1_{macro}$ ($\uparrow$)
OE 0.7083 0.5808 0.6383
CLA 0.5418 0.7143 0.6162
$GAP_{multi}$ 0.7854 0.6837 0.7310
@@
Table 5: Summary statistics of other evaluation measures. Since CLA strictly optimizes for minimizing FNR, it indeed achieves the highest Recall (1 - FNR). However, this comes at the cost of losing out on Precision. Since $GAP_{multi}$ jointly minimizes both TPR and TNR, it performs best both in terms of Precision and F1 scores across three losses.
@@
Table 5 shows summary statistics for Precision, Recall and F1 for multi-label classification setup at the macro level. The CLA approach, characterized by its emphasis on minimizing the False Negative Rate (FNR), inherently yields the highest Recall. However, this optimization strategy comes at the expense of Precision and F1, both being lower than OE. Our proposed $GAP_{multi}$ loss, designed to jointly minimize True Positive Rate (TPR) and True Negative Rate (TNR), emerges as the optimal performing loss function in terms of both Precision and F1 score across the evaluated losses.
8.2 Runtime Performance
@@
Table 6 shows runtime performance results over the training dataset of 36k posts, where we report the average time per epoch, number of epochs till convergence, total runtime, and the extra time ($\Delta$) for losses compared to OE.
Since OE optimizes for weighted Binary Cross Entropy (wBCE), it takes the least time per epoch and the least number of epochs to converge. $GAP_{multi}$ takes all the steps for computing the overall loss in addition to calculating $\prescript{G\mkern-0.5mu}{}{C}_{2}$ pairwise losses and the balanced loss. Hence, $GAP_{multi}$ takes additional compute time for solving its intended optimization. The same argument holds true for CLA, since all of them are regularized single objective variants. Although $GAP_{multi}$ does the $\prescript{G\mkern-0.5mu}{}{C}_{2}$ extra computation, hence the extra runtime ($\Delta$: 1167s), it is not that significant (extra 9s per epoch) w.r.t. to OE, while gaining much more in terms of optimization improvement, because the pairwise error computations are all done in parallel. While OE and $GAP_{multi}$ operate on smooth losses their convergence epoch is relatively fast ($\sim$21, $\sim$27). CLA uses a 1-norm loss (Eq. 8), hence the empirical loss surface is not as smooth as the previous two. As such, it is observable that CLA on average takes more epochs ($\sim$41) to converge with a higher $\Delta$.
@@ 9 Discussion
Multi-group target detection brings various challenges and complexities, particularly in balancing the trade-offs between utility and group fairness, as our findings highlight. Achieving fairness in this setting involves mitigating disparities across diverse demographic groups, which is further complicated by various issues: data imbalance, base rate disparities, and the multi-label nature of the task. In this section we delve into some broader implications of our model formulation and empirical results. Additionally, we discuss its applicability in other fairness-sensitive domains and models, and the opportunities it presents for future work.
@@
Fairness measures must be carefully selected to align with the requirements of the task, context, and stakeholders involved. In target-group detection, where posts may simultaneously target multiple groups, the symmetric error costs motivates a fairness measure that enforces equitable treatment across all groups. In this context, we adopted Accuracy Parity (AP) as the fairness criterion, as it directly addresses disparities in group-level predictive accuracy. AP measures the extent to which no group, particularly those with smaller representation or lower base rates, experiences disproportionate harm. This is accomplished using the $GAP_{multi}$ loss function, which equivalently minimizes for AP. The formulation enforces that fairness is incorporated directly into the model training process, enabling balanced performance without requiring additional pre- or post-processing steps.
@@
Ensuring fairness across a growing number of demographic groups $G$ poses computational challenges, particularly for methods relying on sequential global metrics like deviations from mean. In contrast, our $GAP_{multi}$ formulation leverages a pairwise structure, computing error deviations between all group pairs $\prescript{G\mkern-0.5mu}{}{C}_{2}$ independently. This design enables parallelization, integrating seamlessly with modern deep learning pipelines and exploiting GPU acceleration for efficiency. Despite the quadratic growth in terms of $G$, the $GAP_{multi}$ loss retains its convergence properties, ensuring stable optimization even for high group cardinalities, making it robust for real-world multi-group scenarios.
@@
For the 7-group setting in Table 3, we see that by optimizing for group-related errors alongside overall error, both $GAP_{multi}$ and CLA driven classifiers achieve better overall performance (in terms of Avg. BA), and also achieve their intended group-specific objective. Given our multi-label classification where the groups serve as the labels, we hypothesize that the group indicator in each classification node gives an extra feature dimension. This likely enables the model to learn something more about the data than it could without the group-specific information. By incorporating group-associated terms, both $GAP_{multi}$ and CLA have a modified surface compared to OE, allowing convergence to a better optima. We see this pattern emerging in some of the 2-group setting: for example, the Latinx vs. Middle Eastern group-pair in Appendix C.
Model Multiplicity [4] refers to the phenomenon where different models achieve similar overall utility but produce varying predictions. With target-group detection, classifiers trained using CLA and $GAP_{multi}$ exhibit similar overall accuracy but diverge significantly in their group-level performance. In this scenario, AP as a fairness measure provides valuable insight by highlighting the amount of bias across groups, with $GAP_{multi}$ having the least pairwise performance gap across groups.
Incorporating symmetric error considerations, where type-I and type-II errors are considered equally harmful, not only enhances the fairness of our multi target-group detection model but also opens new avenues to explore similar scenarios and challenges in other domains.
For example, this type of error is associated with tasks of inferring preferences often arising in recommender systems such as Netflix or Spotify. Such systems infer user preferences to suggest movies or songs, which may significantly vary across different demographic groups. A symmetric error approach suggests that recommendations are not skewed toward the preferences of the statistical majority group. By recognizing and addressing the symmetric nature of error costs across different groups with similar intent, we can achieve a balanced perspective on fairness and more equitable outcomes in algorithmic decision-making processes and avoid any portability traps [41].
Fair target-group detection is also relevant in the context of training and deploying Large Language Models (LLMs), such as role prompting LLMs to respond as as a specific character or persona (essentially guiding its tone, style, and reasoning approach) [46, 37, 43]. Studies have found that LLMs contain bias with respect to protected characteristics such as gender and race [34, 27]. Integrating explicit target group detection into LLM training pipelines, particularly through reinforcement learning from human feedback (RLHF), could mitigate bias by dynamically aligning model outputs with fairness constraints across demographic groups.

### v3

@@ 8 Results
Results show that $GAP_{multi}$ not only achieves better group-fairness, but also maintains competitive overall accuracy. This highlights the effectiveness of $GAP_{multi}$ in addressing the unique challenges of multi-group target detection. All reported results in the main text are for the 7-group setting in MHS Corpus and 5-group setting in HateXplain. Additional results are shown in Appendix C. Performance on selected $\prescript{|G|\mkern-0.5mu}{}{C}_{2}$ pairs are in Appendix D.
Table 3: Balanced Accuracy (BA) results on MHS for each loss function (OE and CLA baselines vs. $GAP_{multi}$) for the 7 demographic groups averaged over five runs. We also color which group exhibits the maximum and minimum BA values for each loss over the 7 groups. The difference between this maximum vs. minimum is shown in the Max. Diff. column (for BA, we want to minimize this difference). $GAP_{multi}$ achieves a lower maximum difference (Max. Diff. = 5.54) than either baseline loss functions, evident from the visualization in Fig. 2. $GAP_{multi}$ also achieves the highest (macro) average BA across groups (Avg. BA = 81.97), and the best BA for most of the groups (#Best BA column = 5/7). We also report the standard deviation of Max. Diff and Avg. BA over five runs, with CLA and $GAP_{multi}$ having similar deviation. $GAP_{multi}$ achieves the best fairness performance while also retaining overall utility.
MHS Corpus - Balanced Accuracy (BA) ($\uparrow$)
@@
[…]
Table 3 shows results for the MHS Corpus dataset, with an equivalent visualization is shown in Fig. 2. We show BA values across groups averaged over five test set runs and compare the two baseline losses (OE and CLA) vs. $GAP_{multi}$. Recall that overall error (OE) minimizes $\overline{err}_{overall}$ without any fairness constraints. For the Black group, which is the statistical majority group in the dataset (see App. Fig. 5), optimizing OE loss yields the highest BA (BA is generally higher for this group than others). Performance for the other groups—and in particular for the Native American group, which is a statistical minority—varies starkly across the different optimization strategies. The worst performing group for each loss (i.e., having minimum BA values) is also seen to exhibit the best performance under $GAP_{multi}$. This means that $GAP_{multi}$ yields the best performance according to a min-max criteria [12].
Figure 2: Visualization of the BA values achieved by each loss over the 7 demographic groups in MHS corpus. The maximum difference (Max. Diff.) between the maximum and minimum BA achieved for each loss across groups is also shown. $GAP_{multi}$ performs best with lowest Max Diff. of 5.5, showcasing that it greatly reduces the disparities in performance across groups vs. other losses.
We show the maximum difference (Max. Diff.) between the groups in Table 3, which captures the gap in performance between worst vs. best performing groups. We see that solely optimizing for OE fails to account for variations in group performance, resulting in the highest difference values (Max. Diff. of 21.9). In contrast, both $GAP_{multi}$ and CLA consider group performance alongside OE, leading to substantially lower Max. Diff. in comparison. Notably, $GAP_{multi}$ optimizes for balanced error rates across groups, exhibiting the smallest difference (Max. Diff. of 5.5). Table 3 reports average BA (macro) obtained by each loss, with $GAP_{multi}$ performing best (Avg. BA: 81.97). We hypothesize that by incorporating group information in the loss during model training, both group-informed losses (CLA and $GAP_{multi}$) find better local optima vs. OE. The “#Best BA” column reports for each loss function how many groups achieved the best BA for that loss. We see $GAP_{multi}$ performs best for five groups, while OE is best for the other 2 groups (Black, Middle Eastern). This further shows that $GAP_{multi}$ does not prioritize the performance of one group over others during optimization, thereby being the best performing loss for most of the groups.
Table 4: Balanced Accuracy (BA) results for each loss function (OE and CLA baselines vs. $GAP_{multi}$) for the 5 demographic groups on test data for HateXplain. For each loss, we color which group exhibits the maximum and minimum BA values achieved for that loss over the 5 groups, with the difference between this maximum vs. minimum shown in the Max. Diff. column. $GAP_{multi}$ achieves a lower maximum difference (Max. Diff. = 5.19) than either baseline loss functions. We also report the standard deviation of Max. Diff and Avg. BA over five runs. $GAP_{multi}$ achieves the best fairness performance while also retaining overall utility.
HateXplain - Balanced Accuracy (BA) ($\uparrow$)
Loss African Arab Asian Caucasian Hispanic Max. Diff. ($\downarrow$) Avg. BA ($\uparrow$) #Best BA ($\uparrow$)
[…]
For HateXplain dataset, the results presented in Table 4 show a consistent trend to our findings. The same experimental setup reveals that $GAP_{multi}$ again achieves the most balanced group performance, with the smallest Max. Diff. (5.19) between groups and similar overall utility (74.66). For this dataset, CLA achieves the highest Avg. BA (74.84), but both OE and $GAP_{multi}$ achieve statistically comparable performance. Meanwhile, $GAP_{multi}$ achieves a statistically significant reduction in Max. Diff.. The consistent strong performance of the proposed approach across two distinct datasets with varying platforms, discourse style, and demographic composition shows that $GAP_{multi}$’s optimization for balanced group error is generalizable and not an artifact of a single data source. The comparative reduction in disparity across groups, alongside similar Avg. BA, shows the method’s effectiveness as a fair and utility-preserving loss function for multi target-group identification task. The equivalent visualization is shown in App Fig. 9.
Figure 3: Heatmap of pairwise absolute difference of BA across groups in test set (MHS corpus) as an indicator for bias and disparate impact. OE has the highest performance gap (Max Diff = 21.9) across groups as indicated by the extremes of color, not only across one group- pair but consistently across multiple group pairs. $GAP_{multi}$ has the least spread in pairwise error values (Max. Diff. = 5.5), evident from the flatness of color, indicating least disparate impact across groups.
To further illustrate performance disparities between groups, Figure 3 displays pairwise absolute differences of Balanced Accuracy (BA) across various groups ($|\textit{left - bottom}|$) in the MHS corpus. Higher heatmap values indicate classifier bias, revealing consistent inequities, particularly for overall error (OE) rates, where the (Black, Native American) pair has the largest Max. Diff. (21.9). Two key observations can be drawn. First, the statistical majority Black group dominates performance over the statistical minorities Native American group. Second, disparities persist even for group pairs with similar prevalence (e.g., Middle Eastern vs. White, BA gap: 9.9) because one group may simply be easier to identify than the other. While CLA (optimized for FNR) improves over OE, its misaligned objective (minimizing false negatives, rather than equitable performance), results in a heatmap in-between that of OE and $GAP_{multi}$. In contrast, the $GAP_{multi}$ loss, explicitly designed for balanced group performance, shows substantially fewer extremes (Max. Diff.: 5.5), with smoother heatmap transitions indicating equitable classifier outcomes. Note that the extreme Max. Diff. values presented in Table 3 are reflected as extreme values in these heatmaps for specific group pairs. Similar trends can be observed for the pairwise absolute differences of BA heatmaps on HateXplain (App. Fig. 9), where $GAP_{multi}$ has smoother heatmap transitions compared to the other two losses.
@@
Table 5: Summary statistics of other evaluation measures for the MHS corpus and HateXplain. Since CLA strictly optimizes for minimizing FNR, it indeed achieves the highest Recall (1 - FNR). However, this comes at the cost of losing out on Precision. Since $GAP_{multi}$ maximizes the average of TPR and TNR, it performs best both in terms of Precision and F1 scores across three losses.
MHS Corpus HateXplain
Loss $Prc_{macro}$ ($\uparrow$) $Rec_{macro}$ ($\uparrow$) $F1_{macro}$ ($\uparrow$) $Prc_{macro}$ ($\uparrow$) $Rec_{macro}$ ($\uparrow$) $F1_{macro}$ ($\uparrow$)
OE 0.7083 0.5808 0.6383 0.7379 0.7148 0.7261
CLA 0.5418 0.7143 0.6162 0.5789 0.7492 0.6531
$GAP_{multi}$ 0.7854 0.6837 0.7310 0.7519 0.7365 0.7441
@@
Table 5 shows summary statistics for Precision, Recall and F1 for multi-label classification setup at the macro level on the 7-group MHS corpus and 5-group HateXplain dataset. The CLA approach, characterized by its emphasis on minimizing the False Negative Rate (FNR), yields the highest Recall. However, this optimization strategy comes at the expense of Precision and F1, both being lower than OE across the two datasets. Our proposed $GAP_{multi}$ loss, designed to jointly maximize the average of True Positive Rate (TPR) and True Negative Rate (TNR), is seen to be the best performing loss function in terms of both Precision and F1 score across the evaluated losses.
Table 6 reports Hamming Loss (HL), the average fraction of misclassified labels for the MHS corpus and HateXplain. While CLA (HL: 12.10, 7.22) prioritizes minimizing FNR (which inherently involves asymmetry), at times it disproportionately optimizes for this aspect at the expense of other performance metrics. Consequently, CLA may exhibit poorer HL compared to OE (HL: 7.65, 6.47). In contrast, $GAP_{multi}$ maintains symmetry, achieving a balanced trade-off while maximizing the average of TPR and TNR. We see $GAP_{multi}$ consistently achieves the lowest HL: 6.85, 5.89.
Table 6: HL values across losses for test data in the MHS corpus and HateXplain. Lower values are better. $GAP_{multi}$ optimizes for the average of TPR and TNR, thereby achieving lowest values, indicating better classifier performance.
MHS Corpus HateXplain
OE CLA $GAP_{multi}$ OE CLA $GAP_{multi}$
Hamming Loss (HL) % ($\downarrow$) 7.65 12.10 6.85 6.47 7.22 5.89
Runtime Performance in Table 7 shows results over the training dataset of 36k posts in MHS corpus, where we report the average time per epoch, number of epochs till convergence, total runtime, and the extra time ($\Delta$) for losses compared to OE. Since OE optimizes for weighted Binary Cross Entropy (wBCE), it takes the least time per epoch and the least number of epochs to converge. $GAP_{multi}$ takes all the steps for computing the overall loss in addition to calculating $\prescript{|G|\mkern-0.5mu}{}{C}_{2}$ pairwise losses and the balanced loss. Hence, $GAP_{multi}$ takes additional compute time for solving its intended optimization. The same argument holds true for CLA, since all of them are regularized single objective variants. Although $GAP_{multi}$ does the $\prescript{|G|\mkern-0.5mu}{}{C}_{2}$ extra computation, hence the extra runtime ($\Delta$: 1167s), it is not that significant (extra 9s per epoch) w.r.t. to OE, while gaining much more in terms of optimization improvement, because the pairwise error computations are all done in parallel. While OE and $GAP_{multi}$ operate on smooth losses their convergence epoch is relatively fast ($\sim$21, $\sim$27). CLA uses a 1-norm loss (Eq. 6), hence the empirical loss surface is not as smooth as others. Observably, CLA on average takes more epochs ($\sim$41) to converge with a higher $\Delta$.
@@
A unique advantage of our $GAP_{multi}$ formulation is that pairwise-group errors can be computed in parallel in GPU, thus having $\mathcal{O}(1)$ scaling w.r.t. group cardinality. For illustration, we report runtime per epoch with an increasing number of groups from 2-30 on a synthetic dataset, with each group having 22k samples. Fig. 4 shows that even at #group:30 there is only constant increase in runtime per epoch compared to #group:2 or 6. This is because the calculation of overall error $\overline{err}_{overall}$ is carried out independently and in parallel to the inter-group disparity $dev_{overall}$. The naive version has quadratic scaling resulting in runtime blowup w.r.t. group size, due to the serial computation bottleneck.
Figure 4: Our $GAP_{multi}$ formulation enables parallel GPU computation of pairwise-group errors, achieving constant $\mathcal{O}(1)$ scaling complexity. Both the Optimized and Naive curves are for $GAP_{multi}$ loss with/-out the parallelism involved. The serial bottleneck in the Naive version leads to a quadratic scaling in runtime w.r.t. group size. For our implementation, the Optimized version has runtime per epoch with only constant increase, even when scaling from 2 to 30 demographic groups.
@@ 9 Discussion
Fairness measures must be carefully selected to align with the requirements of the task, context, and stakeholders involved. In multi-group target detection, where posts may simultaneously target multiple groups, the symmetric error costs—symmetric meaning that all groups deserve equal protection—motivate a fairness measure that enforces equitable treatment across all groups. We thus adopted Accuracy Parity (AP) as the fairness criterion, since it directly addresses disparities in group-level predictive accuracy. AP measures the extent to which no group, particularly those with smaller representation or lower base rates, experiences disproportionate harm. This is accomplished using the $GAP_{multi}$ loss function, which equivalently minimizes for AP. The formulation enforces that fairness is incorporated directly into the model training process, enabling balanced performance without requiring additional pre- or post-processing steps.
@@
Ensuring fairness across a growing number of demographic groups $G$ poses computational challenges, particularly for methods relying on sequential global metrics like deviations from mean. In contrast, our $GAP_{multi}$ formulation leverages a pairwise structure, computing error deviations between all group pairs $\prescript{|G|\mkern-0.5mu}{}{C}_{2}$ independently. This design enables parallelization, integrating seamlessly with modern deep learning pipelines and exploiting GPU acceleration for efficiency. Despite the quadratic growth in terms of $|G|$, the $GAP_{multi}$ loss retains its convergence properties, ensuring stable optimization even for high group cardinalities, making it robust for real-world multi-group scenarios.
@@
For the 7-group setting in Table 3, we see that by optimizing for group-related errors alongside overall error, both $GAP_{multi}$ and CLA driven classifiers achieve better overall performance (in terms of Avg. BA) in addition to their intended group-specific objective. Given our multi-label classification where the groups serve as the labels, we hypothesize that the group indicator gives an extra feature dimension. This likely enables the model to learn something more about the data than it could without the group-specific information. By incorporating group-associated indicators, both $GAP_{multi}$ and CLA have a modified surface compared to OE, allowing convergence to a better optima. We see this pattern emerging in some of the 2-group setting: e.g., the Latinx vs. Middle Eastern group-pair in Appendix D.
Crucially, $GAP_{multi}$’s symmetric loss design yields two key advantages over CLA: 1. It avoids imbalance-prone 1-norm penalization, leading to smoother convergence (see Table 7 in Appendix). 2. Its pairwise group-error minimization promotes equitable performance across all labels, which doesn’t risk over-prioritizing certain groups. In the experiments, and aligned with our theoretical proof characterizing the risks of optimizing for EO instead of AP, we find that CLA yields worse performance for statistical minorities, such as Native Americans and Pacific Islanders. Meanwhile, $GAP_{multi}$ minimizes gaps across all groups.
Model Multiplicity [3] refers to the phenomenon wherein several distinct machine learning models achieve nearly identical overall utility, such as accuracy, yet produce divergent predictions for individual data points. This is crucial with target-group detection, where classifiers trained with CLA and $GAP_{multi}$ can exhibit similar overall accuracy but greatly diverge in their group-level performances. This divergence reveals that model selection based solely on overall utility can mask severe disparities in how models treat different demographic subgroups. In this setting, AP serves as a fairness diagnostic providing useful insights into the degree of bias across groups. $GAP_{multi}$ achieves the smallest pairwise performance gap across groups, thereby promoting a more equitable distribution of model utility.